OpenStudy (anonymous):

how do you simplify and what are the steps to: 4x squared over 2x squared-3x-2 x 2x squared-5x+2 over 4x-2

OpenStudy (anonymous):

4x^2? or (4x)^2? They're different

OpenStudy (anonymous):

the first one..

OpenStudy (anonymous):

\[4x ^{2} (2x ^{2}-5x+2)\] *************************** \[( 2x ^{2}-3x-2) (4x-2) \] When you multiply fractions you multiply across Factor your trinomials into 2 binomials and anything you can factor out of the other terms

OpenStudy (anonymous):

\[(2x^2-5x+2) = (2x^2-4x-x+2) = 2x(x-2)-1(x-2) = (2x-1)(x-2)\]

OpenStudy (anonymous):

what do you mean across? as in the denomonators....?

OpenStudy (anonymous):

Yeah numerator multiplies with numerator and same with denominator

OpenStudy (anonymous):

\[(2x^2-3x-2) = (2x^2-4x+x-2) = 2x(x-2)+1(x-2) = (2x+1)(x-2)\]

OpenStudy (anonymous):

okay I think this is gonna help....but its a rational expression, so the answer would need a denominator, right?

OpenStudy (anonymous):

Not necessarily, if everything cancels and the denominator equals 1 then you wouldn't have a denominator.

OpenStudy (anonymous):

okay. I think setting it up in brackets is gonna be easier. thanks for your help :)